OpenStudy (aaronandyson):

the product of the digits of a two digit number is 24.If its unit's digit exceeds twice its tens digit by 2 ;find the number

OpenStudy (aaronandyson):

@UnkleRhaukus

OpenStudy (unklerhaukus):

t * u = 24 u = 2 * t + 2

OpenStudy (aaronandyson):

24/x - 2x/1 = 2?

OpenStudy (aaronandyson):

?